OpenStudy (anonymous):

help please. divide (16x^2-20x+9)/(4x-3) a. 4x-5-3 b.4x-5-9/4x-3 c/4x-2-3/4x-3 d.4x-2+3/4x-3

OpenStudy (anonymous):

@amistre64 could you please check this

OpenStudy (amistre64):

we can do longhand division, or synthetic division ... which are you comfortable with?

OpenStudy (anonymous):

Well the answer I got was C. Synthetic division is ok with me

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i run my sythetic with everything in tact ... the x parts are useless other than being a good reminder in the end: 16x^2 -20x +9, now we divide by a value of x such that 4x-3 = 0: x = 3/4 of course 16x^2 -20x +9 0 12 -6 3/4 16 -8 3 x c R 16x^2 -20x +9 = (4x-3)(4x-2) + 3 which isnt c ...